# Towards Foundation Models for 3D Vision: How Close Are We?

### v1

@@ Abstract
Building a foundation model for 3D vision is a complex challenge that remains unsolved. Towards that goal, it is important to understand the 3D reasoning capabilities of current models as well as identify the gaps between these models and humans. Therefore, we construct a new 3D visual understanding benchmark that covers fundamental 3D vision tasks in the Visual Question Answering (VQA) format. We evaluate state-of-the-art Vision-Language Models (VLMs), specialized models, and human subjects on it. Our results show that VLMs generally perform poorly, while the specialized models are accurate but not robust, failing under geometric perturbations. In contrast, human vision continues to be the most reliable 3D visual system. We further demonstrate that neural networks align more closely with human 3D vision mechanisms compared to classical computer vision methods, and Transformer-based networks such as ViT [17] align more closely with human 3D vision mechanisms than CNNs. We hope our study will benefit the future development of foundation models for 3D vision.
@@ 1 Introduction
3D vision is an important field in computer vision with many fundamental applications, such as navigation, robotics, and augmented/virtual reality. However, foundation models and corresponding benchmarks are largely absent for 3D vision. On one hand, researchers often focus on training specialized models that can solve only a single task, such as depth estimation [37, 65] or optical flow [80]. On the other hand, each task has its own evaluation metrics and there is no benchmark for comparing a single model across different tasks, since the output space of each task is vastly different (e.g., pixel-wise for depth estimation versus $SE(3)$ for camera pose estimation). The pixel-wise dense output required by many tasks also poses challenges in evaluating human performance, making it difficult to study and understand the differences and similarities between models and the human visual system.
@@
To answer the above questions, we construct a new 3D visual understanding benchmark (Sec. 3), which covers fundamental 3D vision tasks including depth estimation, spatial VQA, camera pose estimation, and keypoint matching. While our benchmark is based on existing datasets with accurate ground truth, its key feature is a unified output space across tasks. We formulate all the questions in VQA format so that they can be easily answered by both VLMs and humans, enabling fair comparisons. For example, our depth estimation benchmark asks for binary depth relationships between two pixels rather than a dense depth map. Furthermore, we construct challenging cases with geometric perturbation, such as flipping or rotating the images, to make them diverge from the gravity-aligned views on which the models were trained. Those geometric perturbations are common in applications such as robotics, making it important to test the robustness of models in such scenarios. Compared to existing benchmarks, ours is the first to focus entirely on 3D, as shown in Tab. 1.
We conduct extensive experiments on our benchmark (Sec. 4) by evaluating a variety of models. We test state-of-the-art VLMs including GPT4-Turbo, GPT4-Omni, and Gemini-1.5. We focus on closed-source VLMs because they have the best performance on the 2D leaderboards [6, 51]. We also test the performance of state-of-the-art specialized models on each task. Finally, we record human performance using Amazon Mechanical Turk (MTurk).
@@
Can 2D VLMs solve 3D tasks? No. While VLMs achieve impressive performance on existing VQA benchmarks focusing on 2D, we find that they have poor 3D understanding abilities. None of the existing VLMs achieves human-level performance or performs on par with the specialized model. Surprisingly, on some tasks (e.g., depth estimation), the VLMs perform only marginally better than random guessing, and on geometrically perturbed images, they perform even worse than random guessing.
Are specialized models accurate and robust? They are accurate but not robust. The specialized models have high accuracy in general, and interestingly we find that they are even better than humans on some tasks (e.g., depth estimation and spatial VQA) in the zero-shot setting. However, they are not as robust as humans against geometric perturbations. For example, on depth estimation, the accuracy of MiDaS [65] drops significantly (from 90.4% to 73.9%) on the upside-down images, while human accuracy remains the same (about 83%). Our finding suggests that the specialized models are still vulnerable, despite being trained on a large collection of diverse data [65, 48].
@@
  • We propose a new benchmark with a unified output space for evaluating the 3D understanding capability of existing models.
@@
Table 1: Our benchmark is the only one focusing entirely on 3D vision. It has comparable size (in terms of the number of images) to modern benchmarks for testing VLMs, such as MMBench [51]. Moreover, our benchmark provides human accuracy as a reference for each task and contains challenging samples with geometric perturbation.
@@ 2.1 VLMs and Benchmarks
Several works combine VLMs with 3D understanding. For instance, 3D-LLM [30] and 3D-VisTA [98] train an LLM that can take 3D point clouds as input and complete high-level tasks (e.g., navigation) by leveraging paired 3D-language datasets. In contrast, our paper focuses on benchmarking existing models instead of training. Additionally, we use 2D images as input modality and focus more on low-level 3D vision tasks. Banani et al. [19] trains linear probes on the features of large models to solve 3D tasks such as depth estimation and matching. Compared to them, our method requires no training and can also be applied to closed-sourced VLMs and humans.
Various benchmarks have been proposed to test the scene understanding and visual question answering capability of 2D VLMs [57, 6, 26, 21, 51, 42, 41, 21]. Datasets from a decade ago such as MSCOCO [47] and VQA [1] are large in scale but lack robust and fine-grained evaluation. To resolve this, MMBench [51] proposes an objective evaluation scheme with fine-grained classes of questions. VisIT [6] covers 70 families of instructions and proposes an automatic LLM-based evaluation aligned with human preferences. WildVision [57] generates open-ended questions automatically by leveraging classification datasets. All of these datasets focus on testing the 2D capability of VLMs, whereas our benchmark focuses entirely on 3D.
Concurrent work BLINK [21] collects human performance data on a set of tasks. Some of the tasks are in 3D (e.g., relative depth), while others are in 2D (e.g., IQ test). In contrast to BLINK, our paper focuses entirely on 3D, covering more 3D tasks such as keypoint detection and including more challenging samples with geometric perturbations. Moreover, BLINK evaluates human performance with only 2 subjects (the coauthors), whereas our benchmark evaluates human performance with 162 subjects in depth estimation, 109 subjects in camera pose estimation, 449 subjects in VQA, and 143 subjects in keypoint matching. Our large sample size helps us discover statistically significant differences between models and humans.
A high-level comparison between this work and previous benchmarks is presented in LABEL:{tab:benchmark_comparison}.
@@ 2.3 Specialized 3D Vision Models
Monocular Depth Estimation MiDaS [65] is the first to explore large-scale training on a mixture of datasets. Depth Anything [91] explores semi-supervised learning on unlabeled images. Marigold [37] uses pre-trained diffusion models to generate sharper results. While achieving impressive accuracy, the models are not robust: Zhang et al. [97] shows that a simple strategy can significantly degrade the depth estimation results. Since absolute depth is challenging to predict, DIW [11] explores using humans to annotate the relative depth between two pixels. Our relative depth prediction task is inspired by DIW, but our relative depth ground-truth labels are generated from Lidar sensors rather than from human annotators, making them more reliable. We use MiDaS [65] in this study because it provides both CNN-based and ViT-based variants.
@@
Visual Question Answering (VQA) Tranformer-based VQA methods can be classified into single-stream [12, 44, 96, 45] and two-streams [55, 76, 56, 78], depending on how the images and questions are processed. MDETR [36] explicitly detects the objects and fuses the vision and language stream with a Transformer. We use MDETR as the specialized model for its state-of-the-art accuracy.
@@
Many datasets have been proposed as VQA benchmarks: some focus on semantics [1, 35, 74, 86, 66] while others focus on spatial relationship reasoning [34, 93, 32]. We build our benchmark on top of the widely-used CLEVR [34] dataset.
@@
Camera Pose Estimation This task aims to estimate the 6DoF pose (translation and rotation) of a camera either with respect to a global coordinate system or relative to another camera. Matching-based methods estimate the relative pose using matched keypoints [72, 60, 82, 81, 70, 50, 70, 63], whereas pose regression methods [2, 38, 85, 9, 40, 75, 10] estimate the 6DoF pose directly based on the input images. Our experiments show that none of the models classify pose similar to humans with the highest Cohen’s $\kappa$ score being 0.16.
@@
Keypoint Detection This task involves detecting salient keypoints from an image, typically followed by local feature extraction around those keypoints. Classical computer vision methods such as [54, 28, 68, 69] extract keypoints and features using local information such as image gradients. In contrast, recent deep neural network-based methods such as [15, 4, 84, 18] train CNNs to detect keypoints and extract local features. Our results suggest that humans follow a keypoint detection strategy more similar to neural networks than classical methods.
@@
Keypoint Matching This task involves matching a pixel in one view to the corresponding pixel in another view. Keypoint matching methods can be grouped into two classes: detector-free methods, such as [77, 80, 87, 8], which perform dense matching and avoid the keypoint detection phase; and detector-based methods, which rely on a keypoint detector and local features extracted from those keypoints. Detector-based methods can be further divided into two subclasses: classical methods [54, 28, 68, 69, 53], which use k-Nearest Neighbors in the feature space; and deep learning-based matchers [48, 71, 7], which train neural networks to match the extracted keypoints. Our experiments show that human keypoint matching is more similar to neural network-based methods than classical methods.
3 The 3D Visual Understanding Benchmark
@@
Spatial VQA CLEVR [34] 500 MDETR [36]
@@
We develop a new benchmark for 3D vision tasks based on existing public datasets. Statistics of our benchmark are presented in Tab. 2. The key feature of our benchmark is a unified output space for all sub-tasks, i.e., we form all questions to be multiple-choice so that all models and humans can be easily and fairly compared. In addition to the specialized models, we collect the response from VLMs (GPT4-Turbo, GPT4-Omni, and Gemini-1.5) and humans.
@@ 3.1 Relative Depth Estimation
Dataset We use KITTI [24] for the relative depth estimation analysis. KITTI is a real-world autonomous driving dataset, and we choose it for two reasons: 1) KITTI has rich and accurate annotations, including accurate depth collected by Lidar, as well as semantic segmentation labels. 2) KITTI images contain both natural and man-made objects, providing a high diversity of visual content. Following DIW [11], we sample markers with a 50% probability of being placed either randomly or symmetrically along a horizontal line. We additionally ensure the two pixels are at least 50px apart in space and 1 meter apart in depth to avoid ambiguity. We collect 500 regular images and another 250 geometrically perturbed images by flipping them upside-down.
3.2 Spatial Visual Question Answering
@@ 3.4 Keypoint Matching
Task Definition We give the subject two views of the same scene. We then ask the subject to choose five points in one image and match the corresponding points in the other image.
Dataset We use the Megadepth-1500 dataset [77, 46]. The scenes are historic locations with wide-baseline images. It provides accurate depth and camera pose ground-truth. For the keypoints chosen by the subject, we use the depth and camera pose to find the ground-truth correspondence in the other image, and we run additional forward-backward consistency checks to ensure co-visibility. We randomly sample 450 image pairs and ask the human subject to annotate 5 keypoint pairs on each of them, resulting in 2,250 keypoint pairs.
@@ 4.2 Relative Depth Estimation
Qualitatively, we compare the distribution of the accuracy over different factors. In Fig. 2(a), we compare the accuracy on symmetric and randomly sampled depth point pairs. The performance gap between the two cases is larger for the MiDaS-CNN compared to humans and MiDaS-DPT, showing the stronger alignment between Transformers and humans. Fig. 2(b) shows the accuracy against the difference in depth (meters) for the two-point sampled. The pairs with larger differences in depth are easier to tell apart, so slopes are positive. Comparing the shape of the curve, MiDaS-DPT is more similar to humans than MiDaS-CNN. Finally, in Fig. 2(d), we compare the accuracy of each model on different classes. We merge the original KITTI classes into 5 super-classes. Note the accuracies are normalized by the overall accuracy of each model. The pattern of MiDaS-DPT is the most similar to humans, whereas MiDaS-CNN performs significantly better on traffic signs and worse on buildings and vegetation.
In conclusion, MiDaS-DPT has the error pattern most similar to humans under multiple criteria, suggesting that Transformers are more similar to humans than CNNs in depth perception.
@@
Figure 3: Results on the Spatial VQA benchmark. (a) Our benchmark requires a strong spatial reasoning ability and is very challenging. (b) Even the specialized VQA model MDETR can only achieve 74.4% accuracy. (c) model accuracy drops as the scene complexity grows (more objects). (d) longer questions don’t necessarily lead to worse performance. See text for detailed analysis.
@@
4.3 Spatial Visual Question Answering
@@
Alignment with Humans We compare the accuracy of different models against the scene complexity, measured by the number of objects in the scene. Results are shown in Fig. 3(c). The accuracy of all models drops as the scene complexity increases, while human accuracy is not much affected. We also measure the complexity of the question. While it is difficult to define the exact complexity, we use the number of words in a sentence as a proxy. The results shown in Fig. 3(d) are quite counter-intuitive: the accuracy of all models grow as the question becomes more complex. In contrast, human accuracy drops slightly. None of the models show a strong correlation with humans, highlighting the potentially different ways that models and humans approach the spatial VQA task.
4.4 Relative Camera Pose Estimation

### v2

@@ Abstract
Building a foundation model for 3D vision is a complex challenge that remains unsolved. Towards that goal, it is important to understand the 3D reasoning capabilities of current models as well as identify the gaps between these models and humans. Therefore, we construct a new 3D visual understanding benchmark named UniQA-3D. UniQA-3D covers fundamental 3D vision tasks in the Visual Question Answering (VQA) format. We evaluate state-of-the-art Vision-Language Models (VLMs), specialized models, and human subjects on it. Our results show that VLMs generally perform poorly, while the specialized models are accurate but not robust, failing under geometric perturbations. In contrast, human vision continues to be the most reliable 3D visual system. We further demonstrate that neural networks align more closely with human 3D vision mechanisms compared to classical computer vision methods, and Transformer-based networks such as ViT [17] align more closely with human 3D vision mechanisms than CNNs. We hope our study will benefit the future development of foundation models for 3D vision. Code is available at https://github.com/princeton-vl/UniQA-3D.
@@ 1 Introduction
In comparison, foundation models and benchmarks are largely absent for 3D vision. On one hand, researchers often focus on training specialized models that can solve only a single task, such as depth estimation [37, 65] or optical flow [80]. On the other hand, each task has its own evaluation metrics and there is no benchmark for comparing a single model across different tasks, since the output space of each task is vastly different (e.g., pixel-wise for depth estimation versus $SE(3)$ for camera pose estimation). The pixel-wise dense output required by many tasks also poses challenges in evaluating human performance, making it difficult to study and understand the differences and similarities between models and the human visual system.
@@
To answer the above questions, we construct a new 3D visual understanding benchmark UniQA-3D (UNIfied Visual-Question-Answering for 3D Vision), which covers fundamental 3D vision tasks including depth estimation, spatial reasoning, camera pose estimation, and keypoint matching. While our benchmark is based on existing datasets with accurate ground truth, its key feature is a unified output space across tasks. We formulate all the questions in VQA format so that they can be easily answered by both VLMs and humans, enabling fair comparisons. For example, our depth estimation benchmark asks for binary depth relationships between two pixels rather than a dense depth map. Furthermore, we construct challenging cases with geometric perturbation, such as flipping or rotating the images, to make them diverge from the gravity-aligned views on which the models were trained. Those geometric perturbations are common in applications such as robotics, making it important to test the robustness of models in such scenarios. Compared to existing benchmarks, UniQA-3D is the first to focus entirely on 3D, as shown in Tab. 1.
We conduct extensive experiments on UniQA-3D by evaluating a variety of models (Sec. 4). We test state-of-the-art VLMs including GPT4-Turbo, GPT4-Omni, and Gemini-1.5. We focus on closed-source VLMs because they have the best performance on the 2D leaderboards [6, 51]. We also test the performance of state-of-the-art specialized models on each task. Finally, we record human performance using Amazon Mechanical Turk (MTurk).
@@
Can 2D VLMs solve 3D tasks well? No. While VLMs achieve impressive performance on existing VQA benchmarks focusing on 2D, we find that they have poor 3D understanding abilities. None of the existing VLMs achieves human-level performance or performs on par with the specialized model. Surprisingly, on some tasks (e.g., depth estimation), the VLMs perform only marginally better than random guessing, and on geometrically perturbed images, they perform even worse than random guessing.
Are specialized models accurate and robust? They are accurate but not robust. The specialized models have high accuracy in general, and interestingly we find that they are even better than humans on some tasks (e.g., depth estimation and spatial reasoning) in the zero-shot setting. However, they are not as robust as humans against geometric perturbations. For example, on depth estimation, the accuracy of MiDaS [65] drops significantly (from 90.4% to 73.9%) on the upside-down images, while human accuracy remains the same (about 83%). Our finding suggests that the specialized models are still vulnerable, despite being trained on a large collection of diverse data [65, 48].
@@
  • We propose a new benchmark UniQA-3D with a unified output space for fair comparison of the 3D understanding capability of different models and humans.
@@
Table 1: UniQA-3D has comparable size (in terms of the number of images) to modern benchmarks for testing VLMs, such as MMBench [51]. Moreover, UniQA-3D provides human accuracy and contains challenging samples with geometric perturbation.
@@ 2.1 VLMs and Benchmarks
Several works combine VLMs with 3D understanding. For instance, 3D-LLM [30] and 3D-VisTA [97] train an LLM that can take 3D point clouds as input and complete high-level tasks (e.g., navigation) by leveraging paired 3D-language datasets. In contrast, our paper focuses on benchmarking existing models instead of training. Additionally, we use 2D images as input modality and focus more on low-level 3D vision tasks. Banani et al. [19] trains linear probes on the features of large models to solve 3D tasks such as depth estimation and matching. Compared to them, evaluating on our UniQA-3D requires no training and can be applied to closed-sourced VLMs and humans.
Various benchmarks have been proposed to test the scene understanding and visual question answering capability of 2D VLMs [57, 6, 26, 21, 51, 42, 41, 21]. Datasets from a decade ago such as MSCOCO [47] and VQA [1] are large in scale but lack robust and fine-grained evaluation. To resolve this, MMBench [51] proposes an objective evaluation scheme with fine-grained classes of questions. VisIT [6] covers 70 families of instructions and proposes an automatic LLM-based evaluation aligned with human preferences. WildVision [57] generates open-ended questions automatically by leveraging classification datasets. All of these datasets focus on testing the 2D capability of VLMs, whereas UniQA-3D focuses entirely on 3D.
BLINK [21] evaluates on a diverse set of tasks. Some of the tasks are in 3D (e.g., relative depth), while others are in 2D (e.g., IQ test). In contrast to BLINK, our paper focuses entirely on 3D, covering more 3D tasks such as keypoint detection and including more challenging samples with geometric perturbations. Moreover, BLINK evaluates human performance on only 2 subjects (the coauthors), whereas we evaluate human performance with 162 subjects in depth estimation, 109 subjects in camera pose estimation, 449 subjects in VQA, and 143 subjects in keypoint matching. The large sample size of our UniQA-3D helps us discover statistically significant differences between models and humans.
@@ 2.3 Specialized 3D Vision Models
Monocular Depth Estimation. MiDaS [65] is the first to explore large-scale training on a mixture of datasets, followed by more recent works [91, 37]. Since absolute depth is challenging to predict, DIW [11] explores using humans to annotate the relative depth between two pixels. Our relative depth prediction task is inspired by DIW, but our relative depth ground-truth labels are generated from Lidar sensors rather than from human annotators, making them more reliable. We use MiDaS [65] in this study because it provides both CNN-based and ViT-based variants.
@@
Visual Question Answering (VQA). Tranformer-based VQA methods can be classified into single-stream [12, 44, 96, 45] and two-streams [55, 76, 56, 78], depending on how the images and questions are processed. MDETR [36] explicitly detects the objects and fuses the vision and language stream with a Transformer. We use MDETR as the specialized model for its state-of-the-art accuracy.
@@
Many datasets have been proposed as VQA benchmarks: some focus on semantics [1, 35, 74, 86, 66] while others focus on spatial relationship reasoning [34, 93, 32]. We build our benchmark on top of the CLEVR [34] dataset.
@@
Camera Pose Estimation aims to estimate the 6DoF pose (translation and rotation) of a camera either with respect to a global coordinate system or relative to another camera. Matching-based methods estimate the relative pose using matched keypoints [72, 60, 82, 81, 70, 50, 70, 63], whereas pose regression methods [2, 38, 85, 9, 40, 75, 10] estimate the 6DoF pose directly based on the input images. Our experiments show that none of the models classify pose similar to humans with the highest Cohen’s $\kappa$ score being 0.16.
@@
Keypoint Detection involves detecting salient keypoints from an image, typically followed by local feature extraction around those keypoints. Classical computer vision methods such as [54, 28, 68, 69] extract keypoints and features using local information such as image gradients. In contrast, recent deep neural network-based methods such as [15, 4, 84, 18] train CNNs to detect keypoints and extract local features. Our results suggest that humans follow a keypoint detection strategy more similar to neural networks than classical methods.
@@
Keypoint Matching matches a pixel to the corresponding pixel in another view. Keypoint matching methods can be grouped into two classes: detector-free methods [77, 80, 87, 8], which perform dense matching and avoid the keypoint detection phase; and detector-based methods, which rely on a keypoint detector and local features extracted from those keypoints. Detector-based methods can be further divided into two subclasses: classical methods [54, 28, 68, 69, 53], which use k-Nearest Neighbors in the feature space; and deep learning-based matchers [48, 71, 7], which train neural networks to match the extracted keypoints. Our experiments show that human keypoint matching is more similar to neural network-based methods than classical methods.
3 The UniQA-3D Benchmark
@@
Spatial Reasoning CLEVR [34] 500 MDETR [36]
@@
We develop a new benchmark UniQA-3D for 3D vision tasks based on existing public datasets (Tab. 2). The key feature of our benchmark is a unified output space for all sub-tasks, i.e., we form all questions to be multiple-choice so that all models and humans can be easily and fairly compared. In addition to the specialized models, we collect the response from VLMs (GPT4-Turbo, GPT4-Omni, and Gemini-1.5) and humans.
@@ 3.1 Relative Depth Estimation
Dataset We use KITTI [24] for the relative depth estimation analysis. KITTI is a real-world autonomous driving dataset, and we choose it for two reasons: 1) KITTI has rich and accurate annotations, including accurate depth collected by Lidar, as well as semantic segmentation labels. 2) KITTI images contain both natural and man-made objects, providing a high diversity of visual content. Following DIW [11], we sample markers with a 50% probability of being placed either randomly or symmetrically along a horizontal line. We collect 500 regular images and another 250 geometrically perturbed images by flipping them upside-down.
3.2 Spatial Reasoning
@@ 3.4 Keypoint Matching
Task Definition We give the subject two views of a scene. We then ask the subject to choose five points in one image and match the corresponding points in the other image.
Dataset We use the Megadepth-1500 dataset [77, 46]. The scenes are historic locations with wide-baseline images. It provides accurate depth and camera pose ground-truth. For the keypoints chosen by the subject, we use the depth and camera pose to find the ground-truth correspondence in the other image, and we run additional forward-backward consistency checks to ensure co-visibility. We randomly sample 450 image pairs and ask the human subject to annotate 5 keypoint pairs on each, resulting in 2,250 keypoint pairs.
@@ 4.2 Relative Depth Estimation
Qualitatively, we compare the distribution of the accuracy over different factors. In Fig. 2(a), we compare the accuracy on symmetric and randomly sampled depth point pairs. The performance gap between the two cases is larger for the MiDaS-CNN compared to humans and MiDaS-DPT, showing the stronger alignment between Transformers and humans. Fig. 2(b) shows the accuracy against the difference in depth (meters) for the two-point sampled. The pairs with larger differences in depth are easier to tell apart, so slopes are positive. Comparing the shape of the curve, MiDaS-DPT is more similar to humans than MiDaS-CNN. Finally, in Fig. 2(d), we compare the accuracy of each model on different classes. We merge the original KITTI classes into 5 super-classes. Note the accuracies are normalized by the overall accuracy of each model. The pattern of MiDaS-DPT is the most similar to humans, whereas MiDaS-CNN performs significantly better on traffic signs and worse on buildings and vegetation. In conclusion, MiDaS-DPT has the error pattern most similar to humans under multiple criteria, suggesting that Transformers are more similar to humans than CNNs in depth perception.
@@
Figure 3: Results on the spatial reasoning task. (a) Our benchmark requires a strong spatial reasoning ability and is very challenging. (b) Even the specialized VQA model MDETR can only achieve 74.4% accuracy. (c) model accuracy drops as the scene complexity grows (more objects). (d) longer questions don’t necessarily lead to worse performance. See text for detailed analysis.
@@
4.3 Spatial Reasoning
@@
Alignment with Humans We compare the accuracy of different models against the scene complexity, measured by the number of objects in the scene. Results are shown in Fig. 3(c). The accuracy of all models drops as the scene complexity increases, while human accuracy is not much affected. We also measure the complexity of the question. While it is difficult to define the exact complexity, we use the number of words in a sentence as a proxy. The results shown in Fig. 3(d) are quite counter-intuitive: the accuracy of all models grows as the question becomes more complex. In contrast, human accuracy drops slightly. None of the models show a strong correlation with humans, highlighting the potentially different ways that models and humans approach the spatial reasoning task.